OpenStudy (anonymous):

6x + 2y = 0 4x + y = -1 substution method please help

OpenStudy (anonymous):

in substitution method you find the value of one variable in term of another one from one of the equations and then place the value in 2nd eqn to get the result. in this example, 6x + 2y = 0 => 2y = -6x => y = -3x now placing y = -3x in 2nd eqn 4x + y = -1 => 4x + (-3x) = -1 => 4x -3x = -1 => x = -1 y = -3x = 3